OpenStudy (anonymous):

A six-sided number cube with faces numbered 1,2,3,4,5, and 6 is rolled. What is the probability of rolling a 6 or a number less than 3? A=1/18 B=1/6 C=1/3 D=1/2

OpenStudy (kirbykirby):

You need P(roll a 6 OR number less than 3), which suggests using the addition rule: P(roll 6) + P(number less than 3) Now, you'll notice that you can only get one number for P(number less than 3), which suggests using the addition rule on this itself, since you need to consider all possibilities: P(number is 1) + P(number is 2) Thus, you have P(roll 6) + P(roll 2) + P(roll 3) = 1/6 + 1/6 + 1/6 = 1/2

OpenStudy (kirbykirby):

Thus, you have P(roll 6) + P(roll 2) + P(roll 1) = 1/6 + 1/6 + 1/6 = 1/2** sorry typo on the third P( . )

OpenStudy (anonymous):

Thank you very much! :)